# Landau-Lifshitz damping from Lindbladian dissipation in quantum magnets

### v2

@@ 5 Analytical Derivation of Generalized Landau-Lifshitz Equation
It is interesting that for $\vec{m}\neq m_{\infty}\hat{\vec{m}}$ Eq. 7 extends the usual LL equation (1a) systematically, see also App. D. This can be used to raise simulations from a classical LL level to the quantum mechanical level of local mean fields although non-local entanglement is still neglected.
@@ 6 Conclusions and Outlook
In summary, up to now there was a conceptual gap between the established Lindbladian theory for dissipation in open quantum systems and the phenomenological theory of Landau, Lifshitz and Gilbert for the damping of the temporal behavior of magnetic order. In view of the humongous interest in magnetic dynamics this represented a flaw. So far, the differences of the two approaches were emphasized, for instance the change of entropy in Lindblad treatments, see e.g. Ref. [38], and the conservation of entropy in the assumed quantum analogues of the LL(G) equations [11, 14]. In the present study, we filled the conceptual gap and reconciled the fundamental quantum description in terms of a Lindbladian master equation and the phenomenological classical LL(G) description. In addition, a systematic extension for the evolution of the length of the expectation value of $\vec{S}$ has been found. This underlines that the Lindblad approach, though based on a local mean-field theory, reaches beyond a merely classical description.
Three key points constitute the important ingredients: (i) the Lindbladian relaxation is constantly adapted to the instantaneous Hamilton operator and its energies, i.e., the relaxation always acts such as to minimize the instantaneous energy; (ii) the self-reference in the LL(G) equations is explained not by ad hoc modifications of the Schrödinger equation, but by a local mean-field approximation; (iii) finally the internal energy scales of the quantum system need to be much larger than the external ones, here $|J|\gg|h_{0}|$. Key point (i) requires that the internal bath dynamics is faster than the system’s dynamics. This is a standard assumption in the derivation of the derivation of the Lindblad master equation. These ideas allowed us to derive the Landau-Lifshitz equation analytically from Lindbladian dynamics and to confirm this by numerical calculations. We did not observe a trace of Landau-Lifshitz-Gilbert dynamics for large values of $\lambda=\gamma/J$.
@@
Another extension can consist in using Eq. (7) instead of the LL equation in atomistic spin simulations [9, 42, 43, 10]. Such approaches are also employed to describe systems of Josephson junctions [44, 45, 46]. This would allow for taking changes in the spin length into account, promoting the classical calculation to a local mean-field one if the external fields are weak with respect to the internal ones. If the weak-field limit is not justified one can equally well resort to Eq. (5) which is based on a local mean-field approach, but does not require a particular energy hierarchy between internal and external magnetic energies. For including thermal fluctuations by stochastical noise terms, the conditions for reaching equilibrium need to be verified [43].
Recently, high frequencies in the spin dynamics of ferromagnets have been predicted and experimentally identified which are attributed to spin inertia and a resulting nutation term in the Landau-Lifshitz-Gilbert equations [47, 48, 49, 50]. Mathematically, a second time derivative of the magnetization is included. This poses the question whether such a term can result from a Lindblad formalism by extending the above reasoning. Alternatively, one can attempt to understand the observed signatures by Eqs. (5) or (7) directly
@@ Acknowledgements
This research has been supported by the Deutsche Forschungsgemeinschaft (DFG) in project UH 90-14/1 and by the Stiftung Mercator in project KO-2021-0027
@@ Appendix C Effect of Large Landau-Lifshitz Damping
The figures in the article and the analytical calculation presented above in the previous section indicate that the LL equation (1a) captures the limit of weak external fields of the Lindbladian dynamics best. Yet, the damping value used in the numerical computation, $\lambda=0.2$, is fairly small and any possible effect resulting from the LLG equation (1b) on the dynamics would be only $4\%$, see Eq. (2). Hence, we study a larger values of $\lambda$. Yet we stick to values $\lambda<1$ since it is known that the LL equation (1a) yields unphysical behavior for $\lambda>1$. Moreover, we only need to identify possible quadratic effects which distinguish Eq. (1a) from Eq. (2). Thus, to corroborate the observation that the Lindbladian dynamics reduces to LL dynamics rather than to LLG dynamics we study the case $\lambda=0.5$ in Fig. 2.
@@
Of course, the damping takes place faster by a factor of $2.5$ than in Fig. 1. In summary, however, we find again that the agreement for large $J$ in panel (a) is almost perfect while deviations become discernible for intermediate $J$ in panel (b). But the qualitative behavior remains the same. We do not observe any shift in the oscillations of the $S^{x}$ component relative to the oscillations in the corresponding LL curve. Hence, we do not see any sign of a renormalization of the precession term. From the LLG equation (2), one would have expected a $25\%$ effect. Also the dynamics of the $S^{z}$ remains very much the same: it compares similarly to the LL dynamics as we observed in Fig. 1. Definitely, there is no $25\%$ effect. Hence, the weak-field limit is given by the LL dynamics (1a) rather than by the LLG dynamics (1b).
@@ Appendix D Effect of Finite Temperature on the Dissipation
$T_{1}\coloneqq\frac{1}{2}\sum_{l}\gamma_{l}n_{\text{B}}(\Delta_{l})\Big{\langle}\big{[}[B_{l},A]B^{\dagger}_{l}\big{]}+\big{[}B_{l}[A,B^{\dagger}_{l}]\big{]}\Big{\rangle}$ (D.1)
where $n_{\text{B}}$ is the bosonic occupation number at inverse temperature $\beta$ given by $n_{\text{B}}=(\exp(\beta\Delta_{l})-1)^{-1}$ as usual. We stress that $T_{1}$ includes the effects of the Lindblad operators $B_{l}$ inducing decay as well as the effects of the Lindblad operators $B_{l}^{\dagger}$ inducing absorption processes. Here, $\Delta_{l}$ is the energy increment induced by the Lindblad operator $B_{l}$. Specifying to the considered case of a spin in a magnetic field, we only have one $l$ and $B=S^{-}$ with $\Delta=J|\vec{m}|$. Carrying out the calculation we find the extended version of Eq. (4)

### v3

@@ 6 Relaxation of the Length of the Order Parameter
Here, we do not refer to the spin length $S$ determining $\langle\vec{S}^{2}\rangle=S(S+1)$, but to the length of $\vec{m}\coloneqq\langle\vec{S}\rangle$ which we denote by $|\vec{m}|$ or $|m|$ for brevity. It is interesting that beyond the saturation limit $\vec{m}\neq m_{\infty}\hat{\vec{m}}$ the Eq. 9 extends the usual LL equation (1a) systematically, see also App. D. This can be used to raise simulations from a classical LL level to the level of local mean fields although non-local entanglement is still neglected. Note that this implies that local quantum aspects are included, namely the possibility to deal with an order parameter $\vec{m}$ being a spin expectation value which has not a fixed length.
Figure 2: Solutions (solid lines) of Lindbladian (5) with the mean-field self-consistency (8) for the initial length of $\vec{m}$ given in the panels; the initial direction of the magnetization is tilted by $\pi-0.1$ relative to the $z$ direction. The other couplings are $S=1/2$, $J=8h_{0}$, $\lambda=0.2$ so that $\gamma=\lambda J=8/5.$ The dashed curves represent the solution for the same parameters of equation (9) which results in the weak-field limit $h_{0}/J\to 0$ in linear order.
Figure 2 extends the calculations shown in Fig. 1b which started from the saturated value $|\vec{m}|=S$ to smaller values corresponding to an initially smaller value of the order parameter $|\vec{m}|<S$. It becomes evident that the relaxation of the length occurs rapidly in accordance with the larger values of $\gamma=\lambda J$. Interestingly, it appears that first the length of the order parameter enters into saturation before its orientation changes significantly. This is true even if its orientation is antiparallel to the external field, note the initial behavior of $m_{z}$ in Figs. 2a and 2b where the initial order parameter is still non-negligible. In this regime, the approximate leading-order equation (9) still captures the dynamics qualitatively, similar to what we saw in Fig. 1b for $|\vec{m}|=S$. The agreement becomes significantly better if we studies smaller values of $h_{0}/J$.
But Fig. 2c illustrates that the approximate equation fail to reflect the true Lindbladian dynamics, even qualitatively, if the inital order parameter is too small. This was to be expected because the length $|\vec{m}|$ occurs in (9) in the denominator of the relaxational term formed by the double cross product. Here, one may wonder whether the coefficient $C$ in (9) can change sign. This would have important physical consequences because the double cross product with negative sign ensures that the orientation aligns with the external field in the long run $t\to\infty$. A different sign would spoil this wanted relaxational behavior. But from the inequality $|\vec{m}|\leq S$ which is ensured for the quantum mechanical expectation value $\langle\vec{S}\rangle$, it is evident that $C\geq 1$. Still, it cannot be excluded rigorously that the dynamics in (5) or (9) lead to an overshooting $|\vec{m}|\geq S$. But none of our test runs starting from physical reasonable initial conditions showed any sign of such overshooting.
In summary, Eq. (5) using (8) constitutes a justified extension of the LL equation (1a) including the dynamics of the length of the vector valued order parameter.
@@ 7 Conclusions and Outlook
In summary, up to now there was a conceptual gap between the established Lindbladian theory for dissipation in open quantum systems and the phenomenological theory of Landau, Lifshitz, and Gilbert for the damping of the temporal behavior of magnetic order. In view of the humongous interest in magnetic dynamics this represented a caveat. So far, the differences of the two approaches were emphasized, for instance the change of entropy in Lindblad treatments, see e.g. Ref. [38], and the conservation of entropy in the assumed quantum analogues of the LL(G) equations [11, 14]. Other papers succeeded in linking an LL and a Lindblad description for specific models. In the present study, we filled the conceptual gap and reconciled the quantum description in terms of a Lindbladian master equation and the phenomenological classical LL(G) description on a fundamental level. In addition, a systematic extension for the evolution of the length of the expectation value of $\vec{S}$ has been found. This underlines that the Lindblad approach, though based on a local mean-field theory, reaches beyond a merely classical description.
Three key points constitute the important ingredients: (i) the Lindbladian relaxation is constantly adapted to the instantaneous Hamilton operator and its energies, i.e., the relaxation always acts such as to minimize the instantaneous energy; (ii) the self-reference in the LL(G) equations is explained not by ad hoc modifications of the Schrödinger equation, but by a local mean-field approximation; (iii) finally the internal energy scales of the quantum system need to be much larger than the external ones, here $|J|\gg|h_{0}|$. Key point (i) requires that the internal bath dynamics is faster than the system’s dynamics. This is a standard assumption in the derivation of the Lindblad master equation. These ideas allowed us to derive the Landau-Lifshitz equation analytically from Lindbladian dynamics and to confirm this by numerical calculations. We did not observe a trace of Landau-Lifshitz-Gilbert dynamics for larger values of $\lambda=\gamma/J\leq 1$.
@@
Another extension can consist in using Eq. (9) instead of the LL equation in atomistic spin simulations [9, 42, 43, 10]. Such approaches are also employed to describe systems of Josephson junctions [44, 45, 46]. This would allow for taking changes in the spin length into account, promoting the classical calculation to a local mean-field one if the external fields are weak with respect to the internal ones. If the weak-field limit is not justified one can equally well resort to Eq. (5) which is based on a local mean-field approach, but does not require a particular energy hierarchy between internal and external magnetic energies, except that the bath is assumed to be internally fast. For including thermal fluctuations by stochastical noise terms, the conditions for reaching equilibrium need to be verified [43].
Recently, high frequencies in the spin dynamics of ferromagnets have been predicted and experimentally identified which are attributed to spin inertia and a resulting nutation term in the Landau-Lifshitz-Gilbert equations [47, 48, 49, 50]. Mathematically, a second time derivative of the magnetization is included. This poses the question whether such a term can result from a Lindblad formalism by extending the above reasoning.
@@ Acknowledgements
This research has been supported by the Deutsche Forschungsgemeinschaft (DFG) in projects UH 90-14/1 and UH 90-14/2 as well as by the Stiftung Mercator in project KO-2021-0027.
@@ Appendix C Effect of Large Landau-Lifshitz Damping
The figures in the article and the analytical calculation presented above in the previous section indicate that the LL equation (1a) captures the limit of weak external fields of the Lindbladian dynamics best. Yet, the damping value used in the numerical computation, $\lambda=0.2$, is fairly small and any possible effect resulting from the LLG equation (1c) on the dynamics would be only $4\%$, see Eq. (2). Hence, we study a larger values of $\lambda$. Yet we stick to values $\lambda<1$ since it is known that the LL equation (1a) yields unphysical behavior for $\lambda>1$. Moreover, we only need to identify possible quadratic effects which distinguish Eq. (1a) from Eq. (2). Thus, to corroborate the observation that the Lindbladian dynamics reduces to LL dynamics rather than to LLG dynamics we study the case $\lambda=0.5$ in Fig. 3.
@@
Of course, the damping takes place faster by a factor of $2.5$ than in Fig. 1. In summary, however, we find again that the agreement for large $J$ in panel (a) is almost perfect while deviations become discernible for intermediate $J$ in panel (b). But the qualitative behavior remains the same. We do not observe any shift in the oscillations of the $S^{x}$ component relative to the oscillations in the corresponding LL curve. Hence, we do not see any sign of a renormalization of the precession term. From the LLG equation (2), one would have expected a $25\%$ effect. Also the dynamics of the $S^{z}$ remains very much the same: it compares similarly to the LL dynamics as we observed in Fig. 1. Definitely, there is no $25\%$ effect. Hence, the weak-field limit is given by the LL dynamics (1a) rather than by the LLG dynamics (1c).
@@ Appendix D Effect of Finite Temperature on the Dissipation
$T_{1}\coloneqq\frac{1}{2}\sum_{l}\gamma_{l}n_{\text{B}}(\hbar\omega_{l})\Big{\langle}\big{[}[B_{l},A]B^{\dagger}_{l}\big{]}+\big{[}B_{l}[A,B^{\dagger}_{l}]\big{]}\Big{\rangle}$ (D.1)
where $n_{\text{B}}$ is the bosonic occupation number at inverse temperature $\beta$ given by $n_{\text{B}}=(\exp(\beta\hbar\omega_{l})-1)^{-1}$ as usual. We stress that $T_{1}$ includes the effects of the Lindblad operators $B_{l}$ inducing decay as well as the effects of the Lindblad operators $B_{l}^{\dagger}$ inducing absorption processes. Here, $\hbar\omega_{l}$ is the energy increment induced by the Lindblad operator $B_{l}$. Specifying to the considered case of a spin in a magnetic field, we only have one $l=1$ and $B_{1}=S^{-}$ with $\hbar\omega_{1}=J|\vec{m}|$. Carrying out the calculation we find the extended version of Eq. (4)